OpenStudy (anonymous):

graph y=4x-x^3. Using integration find the area bound by the curve and the x axis. THANK YOU

OpenStudy (anonymous):

So the first thing you do when trying to find the area bounded by a curve is to find the intersections, in this case with the x axis.

OpenStudy (anonymous):

i think they are 0 -2 and 0 2?

OpenStudy (anonymous):

oops you're right

myininaya (myininaya):

\[\frac{4x^{1+1}}{2}-\frac{x^{3+1}}{3+1}=2x^2-\frac{x^4}{4}\] you will use this

OpenStudy (anonymous):

yes you're right the intersections are 0,2 and 0,-2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now you have the bounds so you integrate from -2 to 2

OpenStudy (anonymous):

wrong

myininaya (myininaya):

\[4x-x^3=x(4-x^2)=x(2-x)(2+x)=0 =>x=0,x=2,x=-2\]

OpenStudy (anonymous):

so i just plug in 2 into the formula then plug in -2 and subtract those two answers?

OpenStudy (anonymous):

the integral will be zero , because its an odd function

OpenStudy (anonymous):

but the area wont be zero

OpenStudy (anonymous):

the answer that my book says is 0..

OpenStudy (anonymous):

yeh well they are wrong , if they asked for area

OpenStudy (anonymous):

\[\int\limits_{-2}^{2}4x-x^3dx\] is equal to \[2x^2-1/4x^4 from -2 \to 2\] is equal to 4-4 which would equal 0

OpenStudy (amistre64):

i still get turned around on what they really want in thos

OpenStudy (anonymous):

you need to find the integral from 0 to 2

OpenStudy (anonymous):

however elecengineer is correct

OpenStudy (anonymous):

then double it

myininaya (myininaya):

\[\int\limits_{-2}^{0}f(x)dx+ \int\limits_{0}^{2}f(x)dx\]

OpenStudy (anonymous):

there is an area of 4 before x=0 and an area of 4 after x=0 which =8

OpenStudy (anonymous):

^ still one of them needs absolute value, the first one

OpenStudy (anonymous):

so the answer is 0 or not?

OpenStudy (anonymous):

if they asked for "integral" then yes

OpenStudy (anonymous):

if they asked for "area" then no

myininaya (myininaya):

integral or net area*

OpenStudy (anonymous):

elecengineer is right if they asked for the integral (which they didn't) then the answer is 0 if they asked for the area (which they did) then the answer is 8

OpenStudy (anonymous):

Thanks everyone

OpenStudy (anonymous):

np